angle a is 4 times greater than angle b and angle c is 20 degrees less than angle b...
i have tried every possible whole number and even busted the equation down to as small as 64th increments.we have not yet studied hours,minutes,seconds..
so i figured it was not able to be figured out..could you please help...
thank you..........sherry

Answer by scott8148(6628)

a+b+c=180

"angle a is 4 times greater than angle b" __ a=4b

"angle c is 20 degrees less than angle b" __ c=b-20

substituting __ (4b)+b+(b-20)=180 __ 6b=200 __ b=100/3

substituting __ a=4(100/3) __ a=400/3

substituting __ c=(100/3)-20 __ c=40/3